Case Study, Chapter 45, Antiarrhythmic Agents

A student nurse has a 1-day observational experience in the CCU. As part of the preparation for the experience, the student needs to know about antiarrhythmic agents including lidocaine, propranolol, and diltiazem. (Learning Objectives 2, 3, 4, and 5)

  1. What are the therapeutic actions for lidocaine and propranolol?
  2. What are the indications and pharmacokinetics for diltiazem?
  3. What are the lifespan considerations for the older adult for the use of antiarrhythmic agents?
  4. What are the key nursing implementation considerations for patients receiving antiarrhythmic agents?
Answer #1

A. What are the therapeutic actions for lidocaine and propranolol?

Lidocaine:

            -Decreasing automaticity of the ventricular cells

            -Decreases depolarization

            -Increases ventricular fibrillation threshold

Propranolol:

            -Membrane stabilizing effect

            -Competitively blocks beta-adrenergic receptors in the heart and kidney

            -Decreases the influence of the sympathetic nervous system.

B. What are the indications and pharmacokinetics for diltiazem?

Indications:

Treatment of paroxysmal SVT, A-Fib, and A-Flutter.

Pharmacokinetics:

            -Oral:

                        -onset= 30-60 minutes

                        -Peak= 2-3 hours

                        -Duration = 6-8 hours

            IV:

                        -onset= Immediate

                        -Peak= 2-3 minutes

                        -Duration = Unknown

C. What are the lifespan considerations for the older adult for the use of antiarrhythmic agents?

-Assess the patients cardiac status closely.

-Assess for impaired liver function

-Obtain baseline vitals for patient

-Inspect for distention of the abdomen.

-Assess the patients’ neurological status and level of consciousness

D. What are the key nursing implementation considerations for patients receiving antiarrhythmic agents?

-Ensure that emergency life support equipment is readily available.

-Arrange for periodic monitoring of cardiac rhythm when the patient is receiving long-term therapy

-provide comfort measures

-Titrate the does to the smallest amount needed to achieve control of the arrhythmia

-Administer parenteral forms as ordered only if the oral form is not feasible; expect to switch to the oral form as soon as possible

-Continually monitor cardiac rhythm when initiating or changing dose.

Please rate my answer if it is useful to you :) Thank you.